50 mL of a water sample consumed 28.2 mL of EDTA upon titration. The water was treated with
zeolite method and 50 mL of the treated water consumed 2.3 mL of EDTA solution. 50 mL of
SHW required 21 mL of EDTA. Calculate how much of hardness is removed by the zeolite
process? (SHW = 1 g/L)
1
Expert's answer
2021-10-11T06:15:47-0400

50-28.2= 21.8

50-2.3= 46.7

50-21= 29

= 46.7+21.8+29= 97.5g/L